OpenStudy (anonymous):

Please help me with the steps on solving this please would help a lot.. Which equation has no solution? 8 – (5v + 3) = 5v – 5 3m – 6 = 5m + 7 – m 3w + 4 – w = 5w – 2(w – 2) 7y + 9 = 7y – 6 ... not sure on how i would figure this out my minds on like math block. showing the steps would help a great deal thanks

OpenStudy (tyteen4a03):

If any of the equation can be formed into an equation where the left hand side is the same as the right hand side, then it has no real solution. For example, 5x - 4 = 5x - 4 has no real solution.

OpenStudy (anonymous):

I beleive its the first one has no real solution

OpenStudy (anonymous):

not sure how u figure that out ... i mean ik what no soultion means but idk how to find that out

OpenStudy (anonymous):

tyteen is right - so your steps would be to simplify each expression, and then compare. Sue might have overlooked the signs in simplifying...I don't think the first one is the right answer. I get 5-5v = 5v - 5...almost the same, but not. If I continued trying to solve it: -5v=5v-10 -10v = -10 v=-1 Breaking it down: 8-(5v+3) = 5v -5 8-5v-3=5v-1 5-5v=5v-5

OpenStudy (anonymous):

The last one has a solution because we can solve it. Let's try it out.

OpenStudy (anonymous):

So then which one is it I can seem to solve each one... :(

OpenStudy (anonymous):

3m – 6 = 5m + 7 – m This one is m = 5

OpenStudy (anonymous):

Yeah, me too. Any chance you copied one of them incorrectly, Curious?

OpenStudy (anonymous):

8 – (5v + 3) = 5v – 5 3m – 6 = 5m + 7 – m 3w + 4 – w = 5w – 2(w – 2) 7y + 9 = 7y – 6 is what they are werent copied wrong swear

OpenStudy (anonymous):

3w + 4 – w = 5w – 2(w – 2) this one equals 2 = 0

OpenStudy (anonymous):

sorry i meant w = 0

OpenStudy (anonymous):

BTW, first one should be 1, not v=-1.

OpenStudy (anonymous):

Let's look at the last one again: 7y+9 = 7y-6 I think it's this one -- If I try to isolate the variable, I end up with 0 = -15 So maybe we went down the wrong path when we followed tyteen's method. I just skimmed over this one the first time, because the two sides were obviously not the same, but now I've changed my mind.

OpenStudy (anonymous):

ok that makes a bit more sense thanks... :)

OpenStudy (anonymous):

Yea I agree with misthorn! Really well reasoned!

OpenStudy (anonymous):

Teamwork!

OpenStudy (anonymous):

lol yep!